OpenStudy (gitface):

Simplify sin(2arcsin(5/13))

OpenStudy (gitface):

\[\sin (2\sin^{-1} (\frac{ 5 }{ 13 }))\]

OpenStudy (gitface):

Okay I believe I have the answer.  Here's my triangle.